The equation of transverse wave on a rope is
y(x, t) = 7 sin (4.0t − 0.02x)
where y and x are measured in cm and t is expressed in second. Calculate the
maximum speed of a particle on the rope. (5)
****
1
Expert's answer
2020-02-11T10:08:36-0500

maximum speed Vmax =Aw where A - amplitude, w-frequency. From y(x,t)=7sin(4.0t-0.02x) , A=0.07 m , w=4.0 => Vmax=0.07*4=0.28